find the angle of projection at which the horizontal range is twice the maximum height of the projectile
1
Expert's answer
2019-04-08T12:36:45-0400

"a_x=2a_y"

"\tan \alpha = \frac{a_y}{a_x}"

"\tan \alpha = \frac{1}{2}"

"\alpha \approx 26.6 ^\circ"


Answer: "26.6 ^\circ"